Suzanne invested $19,800 in two mutual fund accounts. During the subsequent year, on mutual fund earned 10% and the othe earned 6%. At the end of the year, the total interest earned was $1,492.00. How much did she invest in each fund?

Answer by stanbon(75887)

Suzanne invested $19,800 in two mutual fund accounts. During the subsequent year, on mutual fund earned 10% and the othe earned 6%. At the end of the year, the total interest earned was $1,492.00. How much did she invest in each fund?
-------------------
Equation:
interest + interest = interest
0.10x + 0.06(19,800-x) = 1492
Multiply thru by 100 to get:
10x + 6*19800 - 6x = 149200
4x = 149200 - 6*19800
x = $7600 (amt. invested at 10%)
19800-x = 19800-7600 = $12,200 (amt. invested at 6%)
===============
Cheers,
Stan H.